a printer manufacturing company claims that its new ink-efficient printer can print an average of 1600 pages of word documents with standard deviation of 62. thirty-five of those printers showed a mean of 1585 pages. does this support the company’s claim? use 95% confidence level.


1
Expert's answer
2022-05-15T11:58:00-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1600"

"H_a:\mu\not=1600"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z|> 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{1585-1600}{62/\sqrt{35}}\approx-1.4313"

Since it is observed that "|z|= 1.4313<1.96= z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.4313)=0.152344," and since "p= 0.152344>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 1600, at the "\alpha = 0.05" significance level.